40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 4 20:13:39 2019, gautam, Update, SUS, ETMY suspension oddness ETMY_sensors_1_Feb_2019_2230_PST.pdfETMY_UL.pdfETMY_sensors_timeDomain.pdf
    Reply  Thu Feb 7 19:34:18 2019, gautam, Update, SUS, ETMY suspension oddness ETMY_sensors.pdfpostDiag.pdf
       Reply  Fri Feb 8 20:48:52 2019, gautam, Update, LSC, IFO recovery IMG_7307.JPG
          Reply  Mon Feb 11 15:41:49 2019, gautam, Update, LSC, TRY 60 Hz solved - but clipping persists 
             Reply  Mon Feb 11 16:38:34 2019, gautam, Update, LSC, ETMY OL calibration updated OL_calib_ETMY_PERROR.pdfOL_calib_ETMY_YERROR.pdf
             Reply  Fri Feb 15 18:42:57 2019, gautam, Update, LSC, TRY 60 Hz solved  IMG_7323.JPG
Message ID: 14446     Entry time: Mon Feb 11 15:41:49 2019     In reply to: 14445     Reply to this: 14447   14459
Author: gautam 
Type: Update 
Category: LSC 
Subject: TRY 60 Hz solved - but clipping persists 

Rich came by the 40m to photocopy some pages from Hobbs, and saw me working on the 60 Hz hunting. As I suspected, the problem was being generated in the D040060. This board receives the photodiode signal single-ended, but has a different power ground than the photodiode (even though the PD is plugged into a power strip that claims to come from 1Y4). The mechanism is not entirely clear - the presence of these 60 Hz features seemed to be dependent on the light level on the TRY photodiode (i.e. they were absent when the PSL shutter is closed, and were more prominent when TRY was 0.9 rather than 0.5) but the PD certainly wasn't saturated - the DC signal was only ~100 mV when viewed on a scope. In any case, Rich suggested the simplest test would be to ground the BNC shield bringing TRY to the rack, to the local ground on the board, which I did using a crocodile clip. This did the trick, the TRY signal RMS is now dominated by the ~1 Hz seismic-driven variation.

 On a more pessimistic note - it looks like the elliptical reflector moving did not work, and the clipping in the Y arm persists no. I am able to recover TRY~1 with the yaw offset on the ETM (which is still lower than the 1.06-1.07 Koji reported in Aug 2018, but I can believe that being down to the MC transmission being a few % lower at 15000cts rather than 15500), while the maximum I see without it is ~0.9. This is puzzling, because when the chamber was open, we saw that there was ~1.5" clearance between the edge of the reflector and the beam on an IR card. I suppose the input pointing could have been off by a small amount. So one of the primary vent objectives wasn't acheieved... But I will push ahead with the loss measurement.

ELOG V3.1.3-